# PINGS-X: Physics-Informed Normalized Gaussian Splatting with Axes Alignment for Efficient Super-Resolution of 4D Flow MRI

### v1

@@ Introduction
4D flow MRI is a powerful, non-invasive method for quantifying blood flow velocities (markl20124d) without involving ionizing radiation. It captures time-resolved, three-dimensional velocity fields using phase-contrast MRI, which extends conventional MRI by encoding velocity-sensitive phase shifts into the MR signal. Crucially, the detailed hemodynamic insights from high-resolution 4D flow MRI are vital for predicting the progression of critical conditions such as stenosis and aneurysms (ferdian2023cerebrovascular).
@@
Nevertheless, achieving high spatiotemporal resolution is challenging, as the multi-dimensional acquisition process involving four velocity encodings for each 3D volume and time frame is inherently slow (markl20124d). While modern acceleration techniques like parallel imaging (deshmane2012parallel) and compressed sensing (lustig2007sparse; kontogiannis2022physics) are used, they have inherent limitations. Parallel imaging suffers from noise amplification at high acceleration rates, while compressed sensing relies on sparsity assumptions that fail in complex flow, often blurring structures (gottwald2020pseudo) and underestimating peak velocities (pathrose2021highly). This creates a clinical dilemma: accept rapid, low-resolution scans that risk obscuring information, or perform long, high-fidelity scans that are impractical due to patient discomfort and motion artifacts (zaitsev15motion). This motivates a complementary strategy of post-hoc spatial super-resolution to enhance detail from rapidly acquired data without increasing scan duration.
This potential has motivated various super-resolution strategies for 4D flow MRI (ferdian2023cerebrovascular; saitta2024implicit). Data-driven deep learning methods can learn effective low-to-high-resolution mappings (shit2022srflow; ferdian2023cerebrovascular) but require large patient datasets and often struggle with generalization to unseen cardiovascular domains (ericsson2024generalized). The alternative Physics-Informed Neural Network (PINN) paradigm (raissi2019physics) enforces physical laws on a per-scan basis, circumventing the need for large datasets. However, PINNs must be re-trained for each patient (callmer2025temporal), and their reliance on slow implicit neural representations creates a critical computational bottleneck for high-dimensional 4D flow MRI (saitta2024implicit; callmer2025temporal).
To address these challenges, we draw inspiration from novel view synthesis, where explicit representations have recently surpassed slower, implicit models. Notably, 3D Gaussian splatting (3DGS) (3dgs) achieved state-of-the-art results with training times orders of magnitude faster than neural radiance fields (NeRFs) (nerf). This motivates our key question: Can a bespoke Gaussian-based representation overcome the computational bottlenecks of PINNs for 4D flow MRI super-resolution while retaining theoretical guarantee of convergence?
@@
Each contribution is evaluated on synthetic CFD and real 4D flow MRI datasets with spatially averaged data, creating a more realistic challenging testbed than the sampling‑based settings used in previous work (dirix2022synthesis).
Figure 2: Illustration of the architectural parallel between neural rendering and physics-informed learning. Our work is motivated by the shared use of slow implicit representations (NeRFs and PINNs) and aims to transfer the efficiency of the explicit 3DGS (3dgs) framework to our domain.
@@
First, data-driven deep learning, uses architectures like CNNs to learn direct low-to-high-resolution mappings from paired training data (shit2022srflow; ferdian2023cerebrovascular; long2023super). While powerful, their reliance on large and diverse training datasets is a significant practical challenge, leading to concerns about generalization to unseen cardiovascular domains (ericsson2024generalized).
Second, Physics-Informed Neural Networks (PINNs) offer a compelling alternative by embedding governing laws like the Navier-Stokes equations into the optimization process (raissi2019physics; fathi2020super). This allows them to be trained on a single 4D flow MRI scan without requiring large datasets (fathi2020super; saitta2024implicit). A known synergy exists with neural rendering; for example, the Siren activation function is now often used in PINNs to better represent complex physical fields. Despite this, the fundamental reliance of PINNs on a slow-to-train implicit MLP creates a major computational bottleneck for high-dimensional 4D flow MRI data, limiting their clinical feasibility (saitta2024implicit; callmer2025temporal; cho2023separable).
@@
The computational bottleneck in PINNs mirrors a similar challenge recently overcome in novel view synthesis. Early methods like neural radiance fields (NeRFs) were also limited by slow, implicit MLP-based models (nerf). This limitation was largely solved by 3D Gaussian splatting, which achieved state-of-the-art results with orders-of-magnitude faster training by using an explicit set of 3D Gaussians (3dgs). This successful transition from a slow implicit to a fast explicit representation provides the primary motivation for our work: to bring similar efficiency gains to physics-informed super-resolution.
@@
A few recent studies have explored the intersection of physics and Gaussian representations, but each has key limitations that our work aims to address. Physics-Informed Gaussian Splatting (PIGS) (rensen2024pigs) models time-resolved physical fields by moving 3D Gaussians through the time domain. However, its use of unnormalized Gaussian weights means it lacks theoretical convergence guarantees, which, as our experiments will show, can lead to inferior prediction accuracy. In contrast, Physics-Informed Gaussians (PIG) (kang2024pig) uses a lightweight MLP to process a Gaussian feature embedding. However, by retaining a neural network in its prediction path, it remains an implicit method and, as we found in our experiments, fundamentally shares the same computational overhead as standard PINNs.
@@
Blood flow in large arteries and veins is often approximated as an incompressible Newtonian fluid (lynch2022effects). Subsequently, the velocity field obeys the incompressible Navier–Stokes equation in Eq. (1), which describe the motion of a viscous fluid, and the continuity equation, which describes mass conservation in Eq. (2).
@@
where $z_{i}(\mathbf{x})=\exp\left(-\frac{1}{2}(\mathbf{x}-\bm{\mu}_{i})^{\top}\mathtt{\Sigma}_{i}^{-1}(\mathbf{x}-\bm{\mu}_{i})\right)$ can be viewed as the (unnormalized) influence of $\mathcal{G}_{i}$. This formulation similarly used in PIGS (rensen2024pigs) is simpler than the original 3DGS rendering equation as it is a direct evaluation in the spatiotemporal domain and thus omits the projection and alpha compositing steps. However, the unnormalized sum above has a fundamental limitation: it is not a well-behaved prediction scheme.
@@
NGS can be viewed as transforming the classic (static) Nadaraya-Watson (NW) estimator (nadaraya1964estimating; watson1964smooth) into a dynamic, learnable framework. Unlike the static nature of NW, which uses fixed points and a global bandwidth, our approach optimizes the Gaussian centers ($\bm{\mu}_{i}$) and learns an individual anisotropic covariance/bandwidth ($\mathtt{\Sigma}_{i}$) for each kernel. Coupled with adaptive density control to add or remove Gaussians, this creates a notably more flexible representation for complex physical fields.
@@
Remark. In the case where each bandwidth $h_{ij}$ is of same order of magnitude, i.e. $h_{ij}=O(h)$ $\forall~i,j$, using the bias-variance balancing bandwidth, we attain the best possible convergence rate in the minimax sense (stone1980; stone1982).
@@
Methods
Lid-driven Y-shaped L-shaped
$t$ $\downarrow$
Rel. $\downarrow$
err. (%)
$t$ $\downarrow$
Rel. $\downarrow$
err. (%)
$t$ $\downarrow$
Rel. $\downarrow$
err. (%)
@@ Methods
# Initial
Gaussians
Lid-driven Y-shaped L-shaped
$t$ $\downarrow$
Rel. $\downarrow$
err.(%)
$t$ $\downarrow$
Rel. $\downarrow$
err.(%)
$t$ $\downarrow$
Rel. $\downarrow$
err.(%)
@@
We used a high-resolution 4D Flow MRI dataset of a carotid artery phantom that replicates patient vasculature (ko2019patient), acquired at 0.35 mm spatial and 25 ms temporal resolution. From 34 time frames per cardiac cycle, five frames around the peak systolic phase were used for constructing the 4D dataset. From this data, low-resolution inputs for our $\times 8$ and $\times 64$ super-resolution tasks were generated via spatial averaging. For this task, the model maps 4D spatiotemporal coordinates ($q=4$) to a 4D physical property vector containing the 3D velocity and pressure ($v_{x},v_{y},v_{z},p$), hence $p=4$.
@@
$t$ $\downarrow$
Rel. $\downarrow$
err. (%)
$t$ $\downarrow$
Rel. $\downarrow$
err. (%)
$t$ $\downarrow$
Rel. $\downarrow$
err. (%)
@@
For the task of physics-informed super-resolution, we compared our method, PINGS-X, against several representative baselines: a standard PINN (pinn), PINN with a Siren activation function (siren), XPINN (xpinn) (using both tanh and Siren activations), and the recent Physics-Informed Gaussians (PIG) (kang2024pig), all of which were run for 100,000 epochs. To evaluate against its theoretical foundation, we also included the classic Nadaraya-Watson (NW) estimator as a non-learnable reference. All models were trained using an Adam optimizer. More details are in the supplementary material (D.4).
@@
For synthetic CFD experiments, PINN, Siren, and each XPINN sub-network shared an architecture of 5 hidden layers with 128 neurons and were trained using a full-batch approach. Models with Tanh activation (PINN, XPINN (tanh)) used a learning rate of $10^{-4}$ (wang2023pinnguide), while those with Siren activation (Siren, XPINN (sin)) used $5\times 10^{-6}$ and $\omega_{0}{=}10$. For PIG, we used the official implementation with default settings, while PINGS-X was initialized with 400 Gaussians, trained for 10,000 epochs with a learning rate of $10^{-2}$, and applied splitting/cloning and merging every 100 epochs.
@@
Rel. $\downarrow$
err. (%)
RMSE $\downarrow$
(cm/s)
[…]
Rel. $\downarrow$
err. (%)
RMSE $\downarrow$
(cm/s)
@@ Acknowledgments
This work was partly supported by National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT) (No.RS-2022-NR070832) and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2020-II201373, Artificial Intelligence Graduate School Program(Hanyang University)) and the CoHASS Research Support Grant (No.025637-00001), funded by Nanyang Technological University, Singapore.
PINGS-X: Physics-Informed Normalized Gaussian Splatting with Axes Alignment for Efficient Super-Resolution of 4D Flow MRI
@@ B.2 Proof of Theorem 1
and this implies the convergence in probability (of the $j$th arbitrary component of $\widehat{\mathbf{v}}({\bf{x}})-\mathbf{v}({\bf{x}})$) by Assumption 3. Since the component-wise convergence in probability holds for all $j=1,\ldots,p$, w.l.o.g, the vectorial convergence holds for $\widehat{\mathbf{v}}({\bf{x}})-\mathbf{v}({\bf{x}})$ and so does the convergence with respect to the Euclidean norm, by standard results in probability theory e.g. (Pollard), as required. This completes the proof. $\square$
@@ C.1 Synthetic 2D CFD datasets
We generated synthetic data for incompressible fluid flow using CFD solvers. First, we utilized the commonly used lid-driven cavity flow (abdelmigid2017revisiting), where the lid velocity was set to 1 m/s in a square domain of size 0.05 m, resulting in a Reynolds number of 500. No-slip boundary conditions were applied on the three stationary walls, while the top lid moved at a constant velocity. The pressure gradient normal to all boundaries was set to zero. The simulation was carried out using the open-source software OpenFOAM on a structured grid with $400\times 400$ cells.
@@
Dataset 2D CFD data
4D flow
MRI data
Shape
Lid-driven
Y-shape
L-shape
Carotid
@@ D.5 List of hyperparameters
Sub-networks
2 (Lid-driven),
3 (L-shaped),
4 (Y-shaped)
@@
Sub-networks
2 (Lid-driven),
3 (L-shaped),
4 (Y-shaped)
@@
PIG Configuration
Official
implementation
@@ Appendix F Additional experimental results
Densification
Thres.
Merging
Thres.
Lid driven Y-shape L-shape
$t$ $\downarrow$
Rel. $\downarrow$
err. (%)
$t$ $\downarrow$
Rel. $\downarrow$
err. (%)
$t$ $\downarrow$
Rel. $\downarrow$
err. (%)
@@ F.3 Experimental results on Burger’s equation
$t$  $\downarrow$
Rel. $\downarrow$
err. (%)
$t$  $\downarrow$
Rel. $\downarrow$
err. (%)
$t$  $\downarrow$
Rel. $\downarrow$
err. (%)
Nada.-Watson
1.5 1.24 <<1 2.33 <<1 1.68
@@
$t$  $\downarrow$
Rel. $\downarrow$
err. (%)
$t$  $\downarrow$
Rel. $\downarrow$
err. (%)
$t$  $\downarrow$
Rel. $\downarrow$
err. (%)
Nada.-Watson
<<1 9.51 <<1 10.13 <<1 8.53

### v2

@@ Introduction
4D flow MRI is a powerful, non-invasive method for quantifying blood flow velocities (Markl et al. 2012) without involving ionizing radiation. It captures time-resolved, three-dimensional velocity fields using phase-contrast MRI, which extends conventional MRI by encoding velocity-sensitive phase shifts into the MR signal. Crucially, the detailed hemodynamic insights from high-resolution 4D flow MRI are vital for predicting the progression of critical conditions such as stenosis and aneurysms (Ferdian et al. 2023).
@@
Nevertheless, achieving high spatiotemporal resolution is challenging, as the multi-dimensional acquisition process involving four velocity encodings for each 3D volume and time frame is inherently slow (Markl et al. 2012). While modern acceleration techniques like parallel imaging (Deshmane et al. 2012) and compressed sensing (Lustig et al. 2007; Kontogiannis and Juniper 2022) are used, they have inherent limitations. Parallel imaging suffers from noise amplification at high acceleration rates, while compressed sensing relies on sparsity assumptions that fail in complex flow, often blurring structures (Gottwald et al. 2020) and underestimating peak velocities (Pathrose et al. 2021). This creates a clinical dilemma: accept rapid, low-resolution scans that risk obscuring information, or perform long, high-fidelity scans that are impractical due to patient discomfort and motion artifacts (Zaitsev et al. 2015). This motivates a complementary strategy of post-hoc spatial super-resolution to enhance detail from rapidly acquired data without increasing scan duration.
This potential has motivated various super-resolution strategies for 4D flow MRI (Ferdian et al. 2023; Saitta et al. 2024). Data-driven deep learning methods can learn effective low-to-high-resolution mappings (Shit et al. 2022; Ferdian et al. 2023) but require large patient datasets and often struggle with generalization to unseen cardiovascular domains (Ericsson et al. 2024). The alternative Physics-Informed Neural Network (PINN) paradigm (Raissi et al. 2019b) enforces physical laws on a per-scan basis, circumventing the need for large datasets. However, PINNs must be re-trained for each patient (Callmer et al. 2025), and their reliance on slow implicit neural representations creates a critical computational bottleneck for high-dimensional 4D flow MRI (Saitta et al. 2024; Callmer et al. 2025).
To address these challenges, we draw inspiration from novel view synthesis, where explicit representations have recently surpassed slower, implicit models. Notably, 3D Gaussian splatting (3DGS) (Kerbl et al. 2023) achieved state-of-the-art results with training times orders of magnitude faster than neural radiance fields (NeRFs) (Mildenhall et al. 2020). This motivates our key question: Can a bespoke Gaussian-based representation overcome the computational bottlenecks of PINNs for 4D flow MRI super-resolution while retaining theoretical guarantee of convergence?
@@
Each contribution is evaluated on synthetic CFD and real 4D flow MRI datasets with spatially averaged data, creating a more realistic challenging testbed than the sampling‑based settings used in previous work (Dirix et al. 2022).
Figure 2: Illustration of the architectural parallel between neural rendering and physics-informed learning. Our work is motivated by the shared use of slow implicit representations (NeRFs and PINNs) and aims to transfer the efficiency of the explicit 3DGS (Kerbl et al. 2023) framework to our domain.
@@
First, data-driven deep learning, uses architectures like CNNs to learn direct low-to-high-resolution mappings from paired training data (Shit et al. 2022; Ferdian et al. 2023; Long et al. 2023). While powerful, their reliance on large and diverse training datasets is a significant practical challenge, leading to concerns about generalization to unseen cardiovascular domains (Ericsson et al. 2024).
Second, Physics-Informed Neural Networks (PINNs) offer a compelling alternative by embedding governing laws like the Navier-Stokes equations into the optimization process (Raissi et al. 2019b; Fathi et al. 2020). This allows them to be trained on a single 4D flow MRI scan without requiring large datasets (Fathi et al. 2020; Saitta et al. 2024). A known synergy exists with neural rendering; for example, the Siren activation function is now often used in PINNs to better represent complex physical fields. Despite this, the fundamental reliance of PINNs on a slow-to-train implicit MLP creates a major computational bottleneck for high-dimensional 4D flow MRI data, limiting their clinical feasibility (Saitta et al. 2024; Callmer et al. 2025; Cho et al. 2024).
@@
The computational bottleneck in PINNs mirrors a similar challenge recently overcome in novel view synthesis. Early methods like neural radiance fields (NeRFs) were also limited by slow, implicit MLP-based models (Mildenhall et al. 2020). This limitation was largely solved by 3D Gaussian splatting, which achieved state-of-the-art results with orders-of-magnitude faster training by using an explicit set of 3D Gaussians (Kerbl et al. 2023). This successful transition from a slow implicit to a fast explicit representation provides the primary motivation for our work: to bring similar efficiency gains to physics-informed super-resolution.
@@
A few recent studies have explored the intersection of physics and Gaussian representations, but each has key limitations that our work aims to address. Physics-Informed Gaussian Splatting (PIGS) (Max Rensen and Eisemann 2024) models time-resolved physical fields by moving 3D Gaussians through the time domain. However, its use of unnormalized Gaussian weights means it lacks theoretical convergence guarantees, which, as our experiments will show, can lead to inferior prediction accuracy. In contrast, Physics-Informed Gaussians (PIG) (Kang et al. 2024) uses a lightweight MLP to process a Gaussian feature embedding. However, by retaining a neural network in its prediction path, it remains an implicit method and, as we found in our experiments, fundamentally shares the same computational overhead as standard PINNs.
@@
Blood flow in large arteries and veins is often approximated as an incompressible Newtonian fluid (Lynch et al. 2022). Subsequently, the velocity field obeys the incompressible Navier–Stokes equation in Eq. (1), which describe the motion of a viscous fluid, and the continuity equation, which describes mass conservation in Eq. (2).
@@
where $z_{i}(\mathbf{x})=\exp\left(-\frac{1}{2}(\mathbf{x}-\bm{\mu}_{i})^{\top}\mathtt{\Sigma}_{i}^{-1}(\mathbf{x}-\bm{\mu}_{i})\right)$ can be viewed as the (unnormalized) influence of $\mathcal{G}_{i}$. This formulation similarly used in PIGS (Max Rensen and Eisemann 2024) is simpler than the original 3DGS rendering equation as it is a direct evaluation in the spatiotemporal domain and thus omits the projection and alpha compositing steps. However, the unnormalized sum above has a fundamental limitation: it is not a well-behaved prediction scheme.
@@
NGS can be viewed as transforming the classic (static) Nadaraya-Watson (NW) estimator (Nadaraya 1964; Watson 1964) into a dynamic, learnable framework. Unlike the static nature of NW, which uses fixed points and a global bandwidth, our approach optimizes the Gaussian centers ($\bm{\mu}_{i}$) and learns an individual anisotropic covariance/bandwidth ($\mathtt{\Sigma}_{i}$) for each kernel. Coupled with adaptive density control to add or remove Gaussians, this creates a notably more flexible representation for complex physical fields.
@@
Remark. In the case where each bandwidth $h_{ij}$ is of same order of magnitude, i.e. $h_{ij}=O(h)$ $\forall~i,j$, using the bias-variance balancing bandwidth, we attain the best possible convergence rate in the minimax sense (Stone 1980, 1982).
@@
Methods Lid-driven Y-shaped L-shaped
$t$ $\downarrow$ (min) Rel. $\downarrow$ err. (%) $t$ $\downarrow$ (min) Rel. $\downarrow$ err. (%) $t$ $\downarrow$ (min) Rel. $\downarrow$ err. (%)
@@
# Initial Gaussians Lid-driven Y-shaped L-shaped
$t$ $\downarrow$ (min) Rel. $\downarrow$ err.(%) $t$ $\downarrow$ (min) Rel. $\downarrow$ err.(%) $t$ $\downarrow$ (min) Rel. $\downarrow$ err.(%)
@@
We used a high-resolution 4D Flow MRI dataset of a carotid artery phantom that replicates patient vasculature (Ko et al. 2019), acquired at 0.35 mm spatial and 25 ms temporal resolution. From 34 time frames per cardiac cycle, five frames around the peak systolic phase were used for constructing the 4D dataset. From this data, low-resolution inputs for our $\times 8$ and $\times 64$ super-resolution tasks were generated via spatial averaging. For this task, the model maps 4D spatiotemporal coordinates ($q=4$) to a 4D physical property vector containing the 3D velocity and pressure ($v_{x},v_{y},v_{z},p$), hence $p=4$.
@@
$t$ $\downarrow$ (min) Rel. $\downarrow$ err. (%) $t$ $\downarrow$ (min) Rel. $\downarrow$ err. (%) $t$ $\downarrow$ (hr) Rel. $\downarrow$ err. (%)
@@
For the task of physics-informed super-resolution, we compared our method, PINGS-X, against several representative baselines: a standard PINN (Raissi et al. 2019a), PINN with a Siren activation function (Sitzmann et al. 2020), XPINN (Jagtap and Karniadakis 2021) (using both tanh and Siren activations), and the recent Physics-Informed Gaussians (PIG) (Kang et al. 2024), all of which were run for 100,000 epochs. To evaluate against its theoretical foundation, we also included the classic Nadaraya-Watson (NW) estimator as a non-learnable reference. All models were trained using an Adam optimizer. More details are in the supplementary material (D.4).
@@
For synthetic CFD experiments, PINN, Siren, and each XPINN sub-network shared an architecture of 5 hidden layers with 128 neurons and were trained using a full-batch approach. Models with Tanh activation (PINN, XPINN (tanh)) used a learning rate of $10^{-4}$ (Wang et al. 2023), while those with Siren activation (Siren, XPINN (sin)) used $5\times 10^{-6}$ and $\omega_{0}{=}10$. For PIG, we used the official implementation with default settings, while PINGS-X was initialized with 400 Gaussians, trained for 10,000 epochs with a learning rate of $10^{-2}$, and applied splitting/cloning and merging every 100 epochs.
@@ Acknowledgments
This work was partly supported by National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT) (No.2021R1A2B5B03002103 and No.RS-2022-NR070832), Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2020-II201373, Artificial Intelligence Graduate School Program(Hanyang University)), Hanyang University (HY-202500000003991) and the CoHASS Research Support Grant (No.025637-00001), funded by Nanyang Technological University, Singapore.
@@ B.2 Proof of Theorem 1
and this implies the convergence in probability (of the $j$th arbitrary component of $\widehat{\mathbf{v}}({\bf{x}})-\mathbf{v}({\bf{x}})$) by Assumption 3. Since the component-wise convergence in probability holds for all $j=1,\ldots,p$, w.l.o.g, the vectorial convergence holds for $\widehat{\mathbf{v}}({\bf{x}})-\mathbf{v}({\bf{x}})$ and so does the convergence with respect to the Euclidean norm, by standard results in probability theory e.g. (Pollard 2002), as required. This completes the proof. $\square$
@@ C.1 Synthetic 2D CFD datasets
We generated synthetic data for incompressible fluid flow using CFD solvers. First, we utilized the commonly used lid-driven cavity flow (AbdelMigid et al. 2017), where the lid velocity was set to 1 m/s in a square domain of size 0.05 m, resulting in a Reynolds number of 500. No-slip boundary conditions were applied on the three stationary walls, while the top lid moved at a constant velocity. The pressure gradient normal to all boundaries was set to zero. The simulation was carried out using the open-source software OpenFOAM on a structured grid with $400\times 400$ cells.
@@
Dataset 2D CFD data 4D flow MRI data
Shape Lid-driven Y-shape L-shape Carotid
@@ D.5 List of hyperparameters
Sub-networks 2 (Lid-driven), 3 (L-shaped), 4 (Y-shaped)
@@
Sub-networks 2 (Lid-driven), 3 (L-shaped), 4 (Y-shaped)
@@
PIG Configuration Official implementation
@@ Appendix F Additional experimental results
Densification Thres. Merging Thres. Lid driven Y-shape L-shape
$t$ $\downarrow$ (min) Rel. $\downarrow$ err. (%) $t$ $\downarrow$ (min) Rel. $\downarrow$ err. (%) $t$ $\downarrow$ (min) Rel. $\downarrow$ err. (%)
@@ F.3 Experimental results on Burger’s equation
$t$  $\downarrow$ (min) Rel. $\downarrow$ err. (%) $t$  $\downarrow$ (min) Rel. $\downarrow$ err. (%) $t$  $\downarrow$ (min) Rel. $\downarrow$ err. (%)
Nada.-Watson 1.5 1.24 <<1 2.33 <<1 1.68
@@
$t$  $\downarrow$ (min) Rel. $\downarrow$ err. (%) $t$  $\downarrow$ (min) Rel. $\downarrow$ err. (%) $t$  $\downarrow$ (min) Rel. $\downarrow$ err. (%)
Nada.-Watson <<1 9.51 <<1 10.13 <<1 8.53